Please solve for x
3(x-6)=2(x+2)
I am having for difficulties w/ this equation....the following is what I have come up with can you please help me.
3(x-6)=2(x+2)
But I cant get any further...THank you so much for your help!!!!

Answer by alexander164(3)

3(x-6)=2(x+2)
3x-18=2x+4
3x+-18=2x+4
3x+-2x=4+18
x=22
Check
3(22-6)=2(22+2)
3(16)=2(24)
48=48
Hope this helps :)